OpenStudy (anonymous):

What is the length of the segment PD? 4 cm 6 cm 10 cm 12 cm

OpenStudy (anonymous):

Indeterminate lol

OpenStudy (amistre64):

you have similar triangles to play with

OpenStudy (anonymous):

What is the length of the segment PD?

OpenStudy (amistre64):

what have you deduced so far?

OpenStudy (amistre64):

what property of similar triangles would help you determine an answer?

OpenStudy (experimentx):

find DQ

OpenStudy (amistre64):

side side side is a good response yes :) namely ratio of side to side

OpenStudy (amistre64):

6x = 4; when x = 4/6 18*4/6 = the part of the lower end of it so 18 - 18*4/6 should give us the required answer

OpenStudy (ujjwal):

they are similar triangles.. all three angles are equal.. n by calculations length of PD is 6 cm

OpenStudy (amistre64):

if i seened it right, i get 18-12

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

P=Q (corr. angle, PR//DE) R=E (corr. angle, PR//DE) Q=Q (common angle) Therefore, QPR is similar to QDE PR/DE = PQ/DQ 6/4 = 18/DQ DQ= 12 PD=PQ-DQ=18-12 =6

OpenStudy (anonymous):

It seems that I am a bit slow. Sorry for that.

OpenStudy (amistre64):

its all those big fancy words you try to use, and hitting that enter key takes up valuable time as well ;)

OpenStudy (anonymous):

Thanks, @amis